OpenStudy (anonymous):

Can someone please help me with the beginning of this integral:

OpenStudy (anonymous):

\[\int\limits_{0}^{1} x^3e^{x^3}dx\]

OpenStudy (amistre64):

\[\int_{0}^{1} x^3e^{x^3}dx\] \[\int_{0}^{1} xx^2e^{x^3}dx\] \[\frac{1}{3}\int_{0}^{1} x~(3x^2e^{x^3})dx\] it looks to be a rather simple by parts now

OpenStudy (amistre64):

u = x^3 u^(1/3) = x du/3u^(2/3) = dx \[\int ue^udx\] \[\frac{1}{3}\int u^{-1/3}e^udu\]if you wanna get fancy

OpenStudy (australopithecus):

the problem with that substitution is that wouldn't you be stuck with an integral that is even harder to deal with

OpenStudy (amistre64):

perhaps, and looking at the wolf, it includes a gamma function

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

is it spose to be this level of difficulty? or are you working at a high level to begin with?

OpenStudy (anonymous):

1st year university maths course

OpenStudy (amistre64):

is the problem written correctly? or is it possible: x^2 e^(x^3) ??

OpenStudy (kainui):

Maybe there's a simpler way we're not thinking of, or the problem was copied wrong. Are you sure it's not\[\int\limits_{0}^{1}x^2e^{x^3}dx\] because that would be much better lol

OpenStudy (anonymous):

nope, it's definitely: (x^3) * (e^(x^3))

OpenStudy (kainui):

remnant have you really learned any other techniques than u-substitution? I'm just wondering what kind of toolset you have to play with lol.

OpenStudy (anonymous):

maybe it was meant as a hard problem to work through. thanks for the help though. just gonna leave it for now as it does seem a lot harder than what I'm accustomed to.

OpenStudy (kainui):

Good luck have fun.

OpenStudy (amistre64):

\[e^u=1+\frac{u}{1!}+\frac{u^2}{2!}+\frac{u^3}{3!}+~...\] \[e^{x^3}=1 +\frac{x^3}{1!}+\frac{x^6}{2!}+\frac{x^9}{3!}+~...\] \[x~e^{x^3}=x +\frac{x^4}{1!}+\frac{x^7}{2!}+\frac{x^{10}}{3!}+~...\] integrating that gives me \[\int x~e^{x^3}=\frac{1}{2}x^2 +\frac{1}{5}\frac{x^5}{1!}+\frac{1}{8}\frac{x^8}{2!}+\frac{x^{10}}{3!}+~...\] is one thought i have, thats prolly wrong tho ;)

OpenStudy (amistre64):

at x=0, thats 0, so at x=1 \[\frac{1}{2}+\frac{1}{5}+\frac{1}{16}+\frac{1}{240}+...\]

OpenStudy (amistre64):

11, not 10 :/

OpenStudy (amistre64):

\[\sum_{n=1}^{inf}\frac{1}{(3n-1)n!}\]

OpenStudy (australopithecus):

\[\int\limits_{0}^{1}x^3e^{x^3}dx\] \[\frac{1}{3}\int\limits_{0}^{1}(x)3x^2e^{x^3}dx\] u = x^3 du = 3x^2dx du/3x^2 = dx insert \[\frac{1}{3}\int\limits_{0}^{1}(x)e^{u}du\] then \[\frac{1}{3} \int\limits_{0}^{1}u^{\frac{1}{3}}e^udu\] blah this would be hard to deal with, maybe you can represent it as a series and solve it that way?

OpenStudy (anonymous):

Thanks for all the help :)

OpenStudy (turingtest):

why did we integrate xe^(x^3) ?\[u=x\implies du=dx\]\[dv=3x^2e^{x^3}dx\implies v=e^{x^3}\]\[\int_0^1x^3e^{x^3}dx=\left.\frac13xe^{x^3}\right|_0^1-\frac13\int_0^1e^{x^3}dx=\frac e3-\frac13\int_0^1e^{x^3}dx\]

OpenStudy (amistre64):

\[x~e^{x^3}=x +\frac{x^4}{1!}+\frac{x^7}{2!}+\frac{x^{10}}{3!}+~...\] forgot it was x^3 and not x :) \[x^3~e^{x^3}=x^3 +\frac{x^6}{1!}+\frac{x^9}{2!}+\frac{x^{12}}{3!}+\frac{x^{15}}{4!}+~...\] \[\int x^3~e^{x^3}=\frac{x^4}{4} +\frac{x^7}{7}+\frac{x^{10}}{20}+\frac{x^{13}}{13*3!}+\frac{x^{16}}{16*4!}+~...\] at x=0 same concept, zeros out at x=1 we have\[\sum_{n=1}^{inf}\frac{1}{n!(1+3n)}\] might prove better

OpenStudy (amistre64):

my summation was not matching the integral, and i figured out why. i need to start with n! = 0! and at the moment im starting off at n!= 1! which is throwing my results off.\[\sum_{n=1}^{inf}\frac{1}{(n-1)!(1+3n)}\]that would work out fine, but it looks to crowded for me so lets adjust it.\[\sum_{n=1-1}^{inf}\frac{1}{(n-1+1)!(1+3(n+1))}\] \[\sum_{n=0}^{inf}\frac{1}{n!(4+3n)}\]And the wolf finally agrees that that is a match :) http://www.wolframalpha.com/input/?i=sum+%28n+%3D+0+to+inf%29+1%2F%28n%21%283n%2B4%29%29